Find x if, 2^sin²x + 2^cos²x = 3


1
Expert's answer
2022-06-15T13:00:48-0400

"2^{\sin^2x}+2^{\cos^2x}=3"

"2^{\sin^2x}+2^{1-\sin^2x}=3"

"2^{\sin^2x}+2\cdot\frac{1}{2^{\sin^2x}}=3"


Let "y=2^{\sin^2x}"


"y+\frac{2}{y}=3"

"y^2-3y+2=0"

"y_1=1", "y_2=2"


"2^{\sin^2x}=1"

"\sin^2x=0"

"\sin x=0"

"x=\pi n, n\in\mathbb{Z}"


"2^{\sin^2x}=2"

"\sin^2x=1"

"\sin x=\pm1"

"x=(-1)^n\frac{\pi}{2}+\pi n, n\in\mathbb{Z}"


Answer:

"x=\pi n, n\in\mathbb{Z}"

"x=(-1)^n\frac{\pi}{2}+\pi n, n\in\mathbb{Z}"